What is the real and imaginary component of matbi ctai ) 2.
Answer #1

solution :- Given Complex Expression is. atbi Ctdi a, b, c, d, m, n all are real numbers, To Real out we have werite (Pota) Im/m act blofil iad tibe e2 toist-di) tegi - car o =ا ما 82 ac tbc + [-ad +bell 2 + d² s/m lactha) (c² + d² Tc2+d²) برو + i[beEulers formula. io by xtiy [xtiy) J x²ty² o = tany re 3/m N/m X Treio 7/m immo e 7 m. Ilos (no) ti sino To By De moivers LawN/m m Imaginally component: Q = sinon.o [a= j x² + y2 otant y in/m tan y ту y = (222 (02 Sinn у bc- ad [x = actbd n/m tiene

거 a ² (² + 9²d ² tb2d2 + b²c2 a2 (²+2 +62(c²+d2 (92+52) (ce+d2) mm Q 9 2462 Sin tan actbd (2+02 bc-ad = 7 Imaginary part or cThanks : if you liked explanation then click on thumbs up :)